What is the concentration of O2 in a freshwater stream in equilibrium with air at 25 degree Celsius and 1.0 bar? Express the answer in grams of o2 per kg of solvent. (Henry’s law constant for O2 is 1.3x10-3 mol/kg bar.at 25 degree Celsius.

Grade:12

1 Answers

Forum 4
one year ago
Solution:

By Henry's law,
Solubility ofO2=KH Pgas
χO2 (in air)=21/100=0.21
∴ pO2=Pair×χO2=0.21 bar
∴ Solubility=1.3×10−3mol/kg bar×0.21 bar
Solubility=2.73×10−4mol/kg=2.73×10^−4×32 g/kg
Solubility=8.736×10−3 g/kg